A boy of mass 40kg slides down a chute inclined at 30 degrees to the horizontal. If the chute is smooth and the boy starts from rest, with what velocity does he pass a point 5 m from the starting point?
1
Expert's answer
2020-03-13T10:48:49-0400

"mg\sin\alpha=ma"


"a=g\sin\alpha"


"S=\frac{v^2}{2a} \to v=\sqrt{2gS\sin\alpha}="


"=\sqrt{2\cdot 9.81\cdot 5\cdot\sin30°}=7" m/s. Answer